OpenStudy (ajanijones):

MEdals.!!!!!! @y1066380 Which ordered pairs are solutions to the equation? y = 3x + 1 A. (6, 19) and (0, 0) B. (6, 19) and (0, 1) C. (6, 18) and (0, 1) D. (19, 6) and (0, 1)

OpenStudy (y1066380):

Just try to put the coordinates into the equation..

OpenStudy (y1066380):

try to put (0, 1) into y = 3x + 1 you'll find the ans

OpenStudy (ajanijones):

Im not sure how to do that my answer im thinking is B. what do you think

OpenStudy (y1066380):

(0,1) is one of the coordinates in B. Put x=0, y=1 into y=3x+1 So, 1=3*0+1 1=1 Since 1=1 is true, so the ans is B